# SF-R You Sure? The Conflicting Role of Star Formation Rates in Constraining the Evolution of Milky Way Analogues in Cosmological Simulations

### v1

@@ 1.1 Milky Way analogues
Our Galaxy supplies us with a wealth of information which we are able to retrieve through surveys like Gaia (Gaia Collaboration et al., 2016) and SDSS (York et al., 2000) to help us understand process such as star formation and stellar evolution, distributions of gas and dust, and chemical enrichment of the interstellar medium (ISM); however, MWAs offer additional information and perspectives not available to us in our own Galaxy.
@@ 1.2 Cosmological simulations
Cosmological simulations may resolve some of the aforementioned obstacles in understanding the MW. These are large-scale simulations that follow the growth of the cosmos from the initial quantum fluctuations, through the formation of large-scale structure and galaxies, and evolved to compare to observations of the Universe today (Somerville & Davé, 2015; Vogelsberger et al., 2020) With these Universes-in-a-box, measurements of galaxy properties are simplified to read-outs of specified quantities, offering both consistency and precision not attainable with observations. Further, cosmological simulations provide the opportunity of time-reversal, making it possible to follow a single galaxy’s quasi-continuous evolution through cosmic time (Springel et al., 2005).
@@
A number of different “flavours” of cosmological simulations exist, each employing their own physics and numerics on a variety of scales. Two main traditional contenders for numerical approaches are hydrodynamical models, which simultaneously solve for gravity using -body methods and evolve the coupled gas as a fluid – e.g., EAGLE (Schaye et al., 2015; Crain et al., 2015), Illustris (Nelson et al., 2015; Vogelsberger et al., 2014) & IllustrisTNG (Weinberger et al., 2017; Pillepich et al., 2018a), Mufasa (Davé et al., 2016) & Simba (Davé et al., 2019) – and semi-analytical models, which evolve the dark matter and gas components separately using simplified flow equations describing bulk motions in the simulation box – e.g., Santa Cruz SAM (Somerville et al., 2008, 2015; Porter et al., 2014; Brennan et al., 2017) (see e.g., Somerville & Davé (2015); Vogelsberger et al. (2020) for reviews). A newer contender is semi-empirical models, which are tuned specifically to match a range of observations – e.g., UniverseMachine (Behroozi et al., 2013, 2019). Zoom-in simulations also exist, which evolve the Universe in a cosmological box with low resolution, then “zoom in” on select galaxies and rerun the simulation at much higher resolution to capture important physical processes more accurately than possible with subgrid models – e.g., FIRE-2 (Hopkins et al., 2018). Finally, there are bespoke zoom-in simulations which aim to reproduce the MW exactly for a given set of properties – e.g., (Agertz et al., 2021; Beane, 2024). With many excellent simulations at our disposal, ultimately for this work, we choose to make use of EAGLE and IllustrisTNG, as they are both large-scale hydrodynamical models with comparable but slightly different subgrid physics.
@@ 1.3 Purpose and Outline
This is exactly the aim of this work – use the results of the evolutionary analysis of a group of galaxies to refine the selection criteria for MWAs. We establish this selection refinement for a group of MWAs in cosmological simulations, and subsequently assess its implications on observational studies.
@@ 2.2 The IllustrisTNG project
Our galaxy sample is collected from TNG100111We choose TNG100 as it is closest the TNG simulation in size and resolution to Ref-L100N1504 in EAGLE. (Nelson et al., 2019; Springel et al., 2018; Pillepich et al., 2018b; Nelson et al., 2018; Naiman et al., 2018; Marinacci et al., 2018). The suffix “100” denotes the size of the periodic box of this run, a cube of side cMpc with initially each of collisionless dark matter particles of mass and baryonic particles of mass . TNG is calibrated on observations of present-day galaxy properties and statistics including the GSMF (Baldry et al., 2008, 2012; Bernardi et al., 2013; D’Souza et al., 2015), the stellar mass-halo mass relation (Behroozi et al., 2013; Moster et al., 2013), the stellar mass-stellar size relation (Baldry et al., 2012; Shen et al., 2003), the black hole-galaxy mass relation (Kormendy & Ho, 2013; McConnell & Ma, 2013), and the total gas mass content within the virial radius of massive groups (Giodini et al., 2009; Lovisari et al., 2015). In addition to these constraints, TNG is tuned to reproduce the overall shape of the cosmic star formation rate density at (Behroozi et al., 2013; Oesch et al., 2015).
@@ 2.2.1 Relevant IllustrisTNG subgrid physics
TNG includes subgrid routines to model many physical process in galaxy evolution, including stellar formation, evolution, and feedback; chemical enrichment and primordial and metal-line cooling; the formation, accretion, merging, and multi-mode feedback of supermassive black holes (BHs); and the seeding and amplification of cosmic magnetic fields. Here, we provide a concise summary of the prescriptions for star formation, stellar feedback, and AGN feedback – the physics most relevant to this work.
@@
BHs form in sufficiently massive haloes, accrete gas from neighbouring cells, and are merged if found to be within the accretion or feedback region of another BH. Unlike EAGLE, TNG employs two methods of AGN feedback: a low-accretion, kinetic, “radio” mode and a high-accretion, thermal, “quasar” mode. While the thermal mode remains the same as in the original Illustris project, in which thermal energy is injected into the immediately surrounding gas, TNG implements a new model for the kinetic feedback, which implements a kinetic BH-driven wind model shown to be responsible for quenching intermediate- and high-mass halo galaxies in the simulation (Weinberger et al., 2017).
@@ 3 Defining Milky Way analogues
The definition of a MWA is comprised of two components: first, the chosen properties to resemble the MW – the selection parameters – and second, some metric to measure the similarity of analogues to the MW itself. We will address both these criteria thoroughly and in succession, beginning with identifying and motivating our choice in selection parameters, then outlining the method used to calculate our “Milky Way-ness” metric. Once our group of analogues is identified, we perform a deeper examination into our most MW-like galaxies. Particularly, we investigate some of the key properties in our analogue group at , focusing mainly on how efficiently the simulations are reproducing the MW’s colour, with a more detailed look into other propertis in Appendix A.
@@ 3.2 Selection methods
The first method we consider to select MWAs is a simple box cut, or binary weighting system, using specified intervals on SM and SFR and which we call . The second method calculates a MW-ness metric, which we call , using weights to measure the distribution of galaxies relative to the MW itself. The third and final method is an adjustment to the second method, by applying the weights and dividing by a probability density function (pdf) that accounts for a data skew towards low-SM, high-SFR galaxies. This results in an improved measure of MW-ness, , measuring the distribution of galaxies from the MW compared to the background number density. In addition to MWAs, we also identify SM- and SFR-controlled samples using each method. All three selection methods are explained in detail in the following sections. The primary method utilized in this paper is the adjusted weights, and thus from now on we denote the weights by unless otherwise specified.
@@ 3.2.2 Method 2: weights
These weights now provide us with a metric of MW-ness, which we call , on a continuous scale of 0 to 1, with 1 being the most MW-like and 0 being the least. This scale is clearly seen in Figure 1, as the galaxies more similar to the MW are coloured with a higher . This method also produces a better fit to the Gaussian distributions in SM and SFR than the binary weights, making the method of MWA selection preferable; however, there is still a noticeable skew towards low-SM/high-SFR galaxies in the MWA group.
3.2.3 Method 3: Adjusted weights
To optimize our MW-ness metric so that our MWA group better fits a Gaussian distribution, we must find a correction to our weights that accounts for the skew towards low-SM or high-SFR galaxies. To accomplish this, we calculate a pdf from a Gaussian kernel density estimation (KDE) on two dimensions (SM and SFR), using scipy.stats.gaussian_kde(X).pdf(X) on the SM and SFR data X. We then calculate our new weights by dividing the weights calculated in Equation 5 by this pdf:
@@ 3.2.3 Method 3: Adjusted weights
We now have a method of determining MWAs that includes a metric of MW-ness and that offsets the large skew towards low-SM or high-SFR galaxies. The MWA distribution in Figure 1 now better fits the red Gaussian, and the colouring indicates that galaxies most similar to the MW have values concentrated around the MW.
To maintain a normalized weighting system with corresponding to the MW itself, we include the LNB15 MW values for SM and SFR in our calculation, and scale accordingly. This results in a few EAGLE and TNG galaxies with or , so that it is no longer true that higher weights are more MW-like. The adjusted weights compromise between the MW-ness of an individual galaxy and the number density of background galaxies with similar SM or SFR. This technique reduces bias based on properties of an overwhelming background sample, and thus is useful for population averages. The adjusted is then the method used to determine MWA populations in this paper, and so from now on we denote unless specified otherwise. However, we briefly note that when considering how MW-like an individual galaxy is, rather than a population, the original weights are preferred.
The associated selections based on the adjusted weights are shown in Figure 2
@@ 3.3 Top 10 MWAs
Introducing a metric of MW-ness makes it straightforward to pick out the top 10 most MW-like galaxies in both simulations. Recall that measuring the MW-ness of individual galaxies is best done using the weights. In EAGLE, the top 10 MWAs have a factor between 0.71 and 0.96; in TNG, the factor lies between 0.67 and 0.98. The nearly equivalent effective sample sizes of MWAs in both EAGLE and TNG along with this similar range in for the most MW-like galaxies indicates a similar distribution of MW-ness across both simulations, which can also be seen in the bottom pannel of Figure 2. The top 10 MWA galaxy IDs, s, SMs, and SFRs are summarized in Table 2 for both simulations.
@@
The images from both simulations show majority spiral galaxies as expected, suggesting that the combination of SM and SFR may act somewhat as a proxy for general morphology. The EAGLE MWAs appear “redder” than those from TNG, but this is partially an artifact of the filters applied by the simulations to produce the mock observations. EAGLE uses SDSS effective filters u (3543 Å), g (4770 Å), and r (7625 Å), while TNG uses NIRcam filters f070W, f115W, and f200W.
@@ 3.4 Some MWAs are redder than others
As a simple check of our selection method and to see how the simulation output compares to observation, we analyze some of the present-day properties of our MWA group. We begin here with an investigation into the colour of our sample of MWAs in both simulations, but note that more properties are analyzed in Appendix A. The comparison colour used comes again from LNB15, where they have calculated a value of from a box cut of galaxies with similar SM and SFR to the MW, since these quantities are strongly correlated with a galaxy’s colour and so their colour is expected to match the MW’s as well.
@@
shows the distribution of simulated galaxies in colour vs SM space, with contours indicating the position of MWAs in this space in EAGLE (green) and TNG (orange), where the simulated MWA values are calculated from -weighted averages and standard deviations to be in EAGLE and in TNG. Also overlaid are the LNB15 values of colour and SM for the MW indicated by the red crossbars. For the contours and crossbars, an average between the asymmetric uncertainties on the MW’s value was taken. The graphic for TNG includes dust, while the graphic for EAGLE was created by compiling galaxies from the tables RefL0100N1504_DustyMagnitudes (with dust) and RefL0100N1504_Magnitudes (without dust), as the simulation which includes dust is missing a significant number of galaxies, and so these gaps were filled in with the and SM values from the simulation run without dust.
@@ 4.2 Evolutionary histories of Milky Way analogues
These plots show the average and the spread in the SFR and SM as functions of time of the SM-controlled group (black), the SFR-controlled group (blue), and the MWAs (green, EAGLE; orange, TNG). The solid line indicates the weighted median of the group at each time step, and the shaded region confines the to weighted quantiles, where the weights applied are those calculated from the adjusted distribution in Equation (7). Essentially, we have taken our analogue group that was selected on the MW’s present-day SM and SFR (the MWAs(0)) and the resulting MW-ness distribution at , traced all galaxies back along their mpbs, and re-applied the weights at each time step to identify the MWAs at that snapshot. A similar approach was repeated with the control groups.
@@ 4.2.1 Similarities between EAGLE and TNG evolutionary histories
Beginning with the SMHs (right column of Figure 6), we see that the first similarity across both simulations is how the MWAs follow the path of the SM-controlled sample more closely than the SFR-controlled sample. In EAGLE especially, the mass assembly of the MWAs and the SM-controlled group is nearly identical, while the tracks are still close – although distinct – in TNG. However, in both simulations, the MWAs and the SFR-controlled group share minimal overlap, suggesting that present-day SM is a better indicator of a MWA’s mass assembly than SFR. In other words, galaxies with similar present-day SM to the MW evolve along the same track for their SMHs, while galaxies that evolve to the MW’s present-day SFR follow different tracks and end up with comparatively too-low SM.
Another commonality between the simulations appears in the SFHs (left column of Figure 6), in which the MWAs follow more closely to the SM-controlled group in early times and switch to follow more closely to the SFR-controlled group at late times. This crossover is demonstrated in Figure 6 by the intersection between the medians of the two control samples in the SFH plots, with MWAs evolution tracing the SM-controlled sample before this point, but following the SFR-controlled sample after it. This transition gives the MWAs their characteristic SFH profile. At early times when tracing the SM-controlled group, the MWAs are elevated above the SFR-controlled sample; this early burst in SFR is required to evolve to a higher SM at present day, as seen in the SMHs where the MWAs(0) end their evolution with far more SM than the SFR-controlled sample (the green and orange lines representing MWA evolution end above the blue line representing the SFR-controlled sample). At late times, when following more closely to SFR, the MWAs are elevated above the quenching SM-controlled group.
We note that while both simulations see this crossover, in TNG the point at which the medians of the two control samples intersect – and thus when the MWAs begin following more closely to the SFR-controlled group – happens much sooner and leaves a larger separation between control samples than in EAGLE. This earlier and wider distinction between the control groups is due to the rapid quenching in the SM-controlled group in TNG, which paired with the clearer distinction between the MWAs and the SM-controlled group in the SMHs compared to EAGLE, demonstrates that SFR is a more important selection parameter in TNG than in EAGLE.

### v2

@@ 1.1 Milky Way analogues
Our Galaxy supplies us with a wealth of information which we are able to retrieve through surveys like Gaia (Gaia Collaboration et al., 2016) and SDSS (York et al., 2000) to help us understand processes such as star formation and stellar evolution, distributions of gas and dust, and chemical enrichment of the interstellar medium (ISM); however, MWAs offer additional information and perspectives not available to us in our own Galaxy.
@@ 1.2 Cosmological simulations
Cosmological simulations may resolve some of the aforementioned obstacles in understanding the MW. These are large-scale simulations that follow the growth of the cosmos from the initial quantum fluctuations, through the formation of large-scale structure and galaxies, and evolved to compare to observations of the Universe today (Somerville & Davé, 2015; Vogelsberger et al., 2020). With these Universes-in-a-box, measurements of galaxy properties are simplified to read-outs of specified quantities, offering both consistency and precision not attainable with observations. Further, cosmological simulations provide the opportunity of time-reversal, making it possible to follow a single galaxy’s quasi-continuous evolution through cosmic time (Springel et al., 2005).
@@
A number of different “flavours” of cosmological simulations exist, each employing their own physics and numerics on a variety of scales. Two main traditional contenders for numerical approaches are hydrodynamical models, which simultaneously solve for gravity using -body methods and evolve the coupled gas as a fluid – e.g., EAGLE (Schaye et al., 2015; Crain et al., 2015), Illustris (Nelson et al., 2015; Vogelsberger et al., 2014) & IllustrisTNG (Weinberger et al., 2017; Pillepich et al., 2018a), Mufasa (Davé et al., 2016) & Simba (Davé et al., 2019) – and semi-analytical models, which evolve the dark matter and gas components separately using simplified flow equations describing bulk motions in the simulation box – e.g., Santa Cruz SAM (Somerville et al., 2008, 2015; Porter et al., 2014; Brennan et al., 2017) (see e.g., Somerville & Davé (2015); Vogelsberger et al. (2020) for reviews). A newer contender is semi-empirical models, which are tuned specifically to match a range of observations – e.g., UniverseMachine (Behroozi et al., 2013, 2019). Zoom-in simulations also exist, which evolve the Universe in a cosmological box with low resolution, then “zoom in” on select galaxies and rerun the simulation at much higher resolution to capture important physical processes more accurately than possible with subgrid models – e.g., FIRE-2 (Hopkins et al., 2018). Finally, there are bespoke zoom-in simulations which aim to reproduce the MW exactly for a given set of properties – e.g., Guedes et al. (2011); Wang et al. (2015); Grand et al. (2017); Font et al. (2020); Agertz & Kravtsov (2015); Garrison-Kimmel et al. (2014); Sawala et al. (2016); Libeskind et al. (2020); Semenov et al. (2024b). With many excellent simulations at our disposal, ultimately for this work, we choose to make use of EAGLE and IllustrisTNG, as they are both large-scale hydrodynamical models with comparable but slightly different subgrid physics.
@@ 1.3 Purpose and Outline
This is exactly the aim of this work – use the results of the evolutionary analysis of a group of galaxies to refine the selection criteria for MWAs of a specific type. We establish this selection refinement for a group of MWAs in cosmological simulations, and subsequently assess its implications on observational studies.
@@ 2.2 The IllustrisTNG project
Our galaxy sample is collected from TNG100111We choose TNG100 as it is the closest TNG simulation in size and resolution to Ref-L100N1504 in EAGLE. (Nelson et al., 2019; Springel et al., 2018; Pillepich et al., 2018b; Nelson et al., 2018; Naiman et al., 2018; Marinacci et al., 2018). The suffix “100” denotes the size of the periodic box of this run, a cube of side cMpc with initially each of collisionless dark matter particles of mass and baryonic particles of mass . TNG is calibrated on observations of present-day galaxy properties and statistics including the GSMF (Baldry et al., 2008, 2012; Bernardi et al., 2013; D’Souza et al., 2015), the stellar mass-halo mass relation (Behroozi et al., 2013; Moster et al., 2013), the stellar mass-stellar size relation (Baldry et al., 2012; Shen et al., 2003), the black hole-galaxy mass relation (Kormendy & Ho, 2013; McConnell & Ma, 2013), and the total gas mass content within the virial radius of massive groups (Giodini et al., 2009; Lovisari et al., 2015). In addition to these constraints, TNG is tuned to reproduce the overall shape of the cosmic star formation rate density at (Behroozi et al., 2013; Oesch et al., 2015).
@@ 2.2.1 Relevant IllustrisTNG subgrid physics
TNG includes subgrid routines to model many physical processes in galaxy evolution, including stellar formation, evolution, and feedback; chemical enrichment and primordial and metal-line cooling; the formation, accretion, merging, and multi-mode feedback of supermassive black holes (BHs); and the seeding and amplification of cosmic magnetic fields. Here, we provide a concise summary of the prescriptions for star formation, stellar feedback, and AGN feedback – the physics most relevant to this work.
@@
BHs form in halos of mass , accrete gas from neighbouring cells, and are merged if found to be within the accretion or feedback region of another BH. Unlike EAGLE, TNG employs two methods of AGN feedback: a low-accretion, kinetic, “radio” mode and a high-accretion, thermal, “quasar” mode. While the thermal mode remains the same as in the original Illustris project, in which thermal energy is injected into the immediately surrounding gas, TNG implements a new model for the kinetic feedback, which implements a kinetic BH-driven wind model shown to be responsible for quenching intermediate- and high-mass halo galaxies in the simulation (Weinberger et al., 2017).
@@ 3 Defining Milky Way analogues
The definition of a MWA is comprised of two components: first, the chosen properties to resemble the MW – the selection parameters – and second, some metric to measure the similarity of analogues to the MW itself. We will address both of these criteria thoroughly and in succession, beginning with identifying and motivating our choice in selection parameters for our specific science case, then outlining the method used to calculate our “Milky Way-ness” metric. Once our group of analogues is identified, we perform a deeper examination into our most MW-like galaxies. Particularly, we investigate some of the key properties in our analogue group at , focusing mainly on how efficiently the simulations are reproducing the MW’s colour, with a more detailed look into other properties in Appendix A, including chemical enrichment and bulge-to-total ratio.
@@ 3.1 Selection parameters
We note that this definition of MWA may not be consistent with some of the MW’s other present-day properties. However, since these properties are not relevant to our science case of investigating SFHs, it is not necessary for our analogues to match the MW on these properties and therefore any inconsistencies of this nature do not impact nor lessen our results. Thus, any mention of a MW “analogue” throughout the remainder of this paper specifically references our own definition of MWA – a galaxy similar to the MW in terms of its present-day SM and SFR. Any conclusions we draw should not be extrapolated to MWAs of any other type, defined in their own way to address a different research question.
@@ 3.2 Selection methods
The first method we consider to select MWAs is a simple box cut, or binary weighting system, using specified intervals on SM and SFR and which we call . The second method calculates a MW-ness metric, which we denote , using weights to measure the distribution of galaxies relative to the MW itself. The third and final method is an adjustment to the second method, by applying the weights and dividing by a probability density function (pdf) that accounts for a data skew towards low-SM, low-SFR galaxies. This results in an improved measure of MW-ness, , measuring the distribution of galaxies from the MW compared to the background number density. In addition to MWAs, we also identify SM- and SFR-controlled samples using each method. All three selection methods are explained in detail in the following sections. The primary method utilized in this paper is the third weighting method, the weights adjusted against the background galaxy population, and thus from now on we denote the weights by unless otherwise specified.
@@ 3.2.2 Method 2: Distance weights
We call this method the distance weights, as they are related to how “far away” each individual MWA is from the MW itself in SM-SFR space. These weights now provide us with a metric of MW-ness, which we call , on a continuous scale of 0 to 1, with 1 being the most MW-like and 0 being the least. This scale is clearly seen in Figure 1, as the galaxies more similar to the MW are coloured with a higher . This method also produces a better fit to the Gaussian distributions in SM and SFR than the binary weights, making the method of MWA selection preferable; however, there is still a noticeable skew towards low-SM/low-SFR galaxies in the MWA group.
3.2.3 Method 3: Population weights
To optimize our MW-ness metric so that our MWA group better fits a Gaussian distribution, we must find a correction to our weights that accounts for the skew towards low-SM or low-SFR galaxies. To accomplish this, we calculate a pdf from a Gaussian kernel density estimation (KDE) on two dimensions (SM and SFR), using scipy.stats.gaussian_kde(X).pdf(X) on the SM and SFR data X. We then calculate our new weights by dividing the weights calculated in Equation 5 by this pdf:
@@ 3.2.3 Method 3: Population weights
We now have a method of determining MWAs that includes a metric of MW-ness and that offsets the large skew towards low-SM or low-SFR galaxies. The MWA distribution in Figure 1 now better fits the red Gaussian, and the colouring indicates that galaxies most similar to the MW have values concentrated around the MW.
To maintain a normalized weighting system with corresponding to the MW itself, we include the LNB15 MW values for SM and SFR in our calculation, and scale accordingly. This results in a few EAGLE and TNG galaxies with or , so that it is no longer true that higher weights are more MW-like. This weighting system compromise between the MW-ness of an individual galaxy and the number density of background galaxies with similar SM or SFR. This technique reduces bias based on properties of an overwhelming background sample, and thus is useful for population averages, and so we name this method the population weights. The adjusted is then the method used to determine MWA populations in this paper, and so from now on we denote unless specified otherwise. However, we briefly note that when considering how MW-like an individual galaxy is, rather than a population, the original distance weights are preferred. The distance weights are more effective in this case since they tell exactly how similar any one individual galaxy is to the MW itself in terms of its SM and SFR without accounting for the background population, maintaining that higher corresponds to more MW-like.
The associated selections based on the population weights are shown in Figure 2
@@ 3.3 Top 10 MWAs
Introducing a metric of MW-ness makes it straightforward to pick out the top 10 most MW-like galaxies in both simulations. Recall that measuring the MW-ness of individual galaxies is best done using the distance weights . In EAGLE, the top 10 MWAs have a factor between 0.71 and 0.96; in TNG, the factor lies between 0.67 and 0.98. The nearly equivalent effective sample sizes of MWAs in both EAGLE and TNG along with this similar range in for the most MW-like galaxies indicates a similar distribution of MW-ness across both simulations, which can also be seen in the bottom pannel of Figure 2. The top 10 MWA galaxy IDs, s, SMs, and SFRs are summarized in Table 2 for both simulations.
@@
The images from both simulations show majority spiral galaxies as expected, suggesting that the combination of SM and SFR may act somewhat as a proxy for general morphology. The EAGLE MWAs appear “redder” than those from TNG, but this is partially an artifact of the filters applied by the simulations to produce the mock observations. EAGLE uses SDSS effective filters u (3543 Å), g (4770 Å), and r (7625 Å), while TNG uses NIRcam filters f070W, f115W, and f200W for their mock images.
@@ 3.4 Some MWAs are redder than others
As a simple check of our selection method and to see how the simulation output compares to observation, we analyze some of the present-day properties of our MWA group. We begin here with an investigation into the colour of our sample of MWAs in both simulations, but note that more properties including chemical enrichment and bulge-to-total ratio are analyzed in Appendix A. The comparison colour used comes again from LNB15, where they have calculated a value of from a box cut of galaxies with similar SM and SFR to the MW, since these quantities are strongly correlated with a galaxy’s colour and so their colour is expected to match the MW’s as well.
@@
shows the distribution of simulated galaxies in colour vs SM space, with contours indicating the position of MWAs in this space in EAGLE (green) and TNG (orange), where the simulated MWA values are calculated from -weighted averages and standard deviations to be in EAGLE and in TNG. Also overlaid are the LNB15 values of colour and SM for the MW indicated by the red crossbars. For the contours and crossbars, an average between the asymmetric uncertainties on the MW’s value was taken. The graphic for TNG includes dust, while the graphic for EAGLE was created by compiling galaxies from the tables RefL0100N1504_DustyMagnitudes (with dust) and RefL0100N1504_Magnitudes (without dust), as the simulation which includes dust is missing a significant number of galaxies, and so these gaps were filled in with the and SM values from the simulation run without dust. We note that this substitution of data from RefL0100N1504_Magnitudes where galaxies are missing in RefL0100N1504_DustyMagnitudes has no effect on our results, as majority of the missing galaxies have very low weights.
@@ 4.2 Evolutionary histories of Milky Way analogues
These plots show the average and the spread in the SFR and SM as functions of time of the SM-controlled group (black), the SFR-controlled group (blue), and the MWAs (green, EAGLE; orange, TNG). The solid line indicates the weighted median of the group at each time step, and the shaded region confines the to weighted quantiles, where the weights applied are the population weights calculated from the adjusted distribution in Equation (7). Essentially, we have taken our analogue group that was selected on the MW’s present-day SM and SFR (the MWAs(0)) and the resulting MW-ness distribution at , traced all galaxies back along their mpbs, and re-applied the weights at each time step to identify the MWAs at that snapshot. A similar approach was repeated with the control groups.
@@ 4.2.1 Similarities between EAGLE and TNG evolutionary histories
Beginning with the SMHs (top row of Figure 6), we see that the first similarity across both simulations is how the MWAs follow the path of the SM-controlled sample more closely than the SFR-controlled sample. In EAGLE especially, the mass assembly of the MWAs and the SM-controlled group is nearly identical, while the tracks are still close – although distinct – in TNG. However, in both simulations, the MWAs and the SFR-controlled group share minimal overlap, suggesting that present-day SM is a better indicator of a MWA’s mass assembly than SFR. In other words, galaxies with similar present-day SM to the MW evolve along the same track for their SMHs, while galaxies that evolve to the MW’s present-day SFR follow different tracks and end up with comparatively too-low SM.
Another commonality between the simulations appears in the SFHs (bottom row of Figure 6), in which the MWAs follow more closely to the SM-controlled group in early times and switch to follow more closely to the SFR-controlled group at late times. This crossover is demonstrated in Figure 6 by the intersection between the medians of the two control samples in the SFH plots, with MWAs evolution tracing the SM-controlled sample before this point, but following the SFR-controlled sample after it. This transition gives the MWAs their characteristic SFH profile. At early times when tracing the SM-controlled group, the MWAs are elevated above the SFR-controlled sample; this early burst in SFR is required to evolve to a higher SM at present day, as seen in the SMHs where the MWAs(0) end their evolution with far more SM than the SFR-controlled sample (the green and orange lines representing MWA evolution end above the blue line representing the SFR-controlled sample). At late times, when following more closely to SFR, the MWAs are elevated above the quenching SM-controlled group.
We note that while both simulations see this crossover, in TNG the point at which the medians of the two control samples intersect – and thus when the MWAs begin following more closely to the SFR-controlled group – happens sooner and leaves a larger separation between control samples than in EAGLE. This earlier and wider distinction between the control groups is due to the rapid quenching in the SM-controlled group in TNG, which paired with the clearer distinction between the MWAs and the SM-controlled group in the SMHs compared to EAGLE, demonstrates that SFR is a more important selection parameter in TNG than in EAGLE.